OpenStudy (cinniomon):

I need help with the following problem: In each formula, substitute the given values of hte variables. The nfind the value of the remaining variable. V=pi r^2h; V=128 and r=8

OpenStudy (anonymous):

h = 1.57

OpenStudy (anonymous):

V = pir^2h 128 = pi*8^2*h h =pi*8^2/128 therefore h = 1.57